integrate the following integral
∫ (2x – 1) / (x – 1)^2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{2x-1}{(x-1)^2}dx
I = \int \frac{2x-2+1}{(x-1)^2}dx
I = \int \frac{2(x-1)+1}{(x-1)^2}dx
I = \int [\frac{2}{x-1}+\frac{1}{(x-1)^2}]dx
I = \int \frac{2}{x-1}dx+\int \frac{1}{(x-1)^2}dx
I = 2log(x-1)+\int \frac{1}{(x-1)^2}dx
I = 2log(x-1)-\frac{1}{(x-1)}+c